# Explicit Multimodal Graph Modeling for Human-Object Interaction Detection

### v1

@@ Introduction
The challenge of few-shot learning is particularly pronounced in HOI detection, where it manifests as a significant performance disparity between rare and non-rare classes. As evidenced in Table 1, this long-tail distribution problem is severe in many methods. For instance, PPDM and FGAHOI-S exhibit a performance gap of over 10 mAP between their rare and non-rare classes. Similarly, in Figure 4, the GNN-based method SCG also has a significant difference between rare and non-rare classes. However, a key observation from Table 1 is that methods incorporating VLMs substantially mitigate this disparity. This suggests that the rich prior knowledge in VLMs can effectively alleviate the Rare-Non-rare bias. Our own results, illustrated in Figure 4, support this hypothesis: both our MGNM and ViPLO, which leverage CLIP, demonstrate a much narrower performance gap between rare and non-rare classes. Interestingly, ViPLO achieves a better balance than our method. We hypothesize that its use of an additional ViTPose-L model provides another source of rich prior knowledge, playing a similar role like VLMs. Moreover, as shown in Table 1 and Figure 4, an unexpected trend emerges for methods equipped with the $\mathcal{H}$-DETR object detector, these models exhibit a remarkable reversal, where performance on rare classes actually exceeds that on non-rare classes. We posit that this is likely attributable to the unique architectural design of $\mathcal{H}$-DETR, and a deeper investigation into this phenomenon presents a compelling direction for future research. In summary, our analysis indicates that leveraging VLMs and advanced object detectors represents a highly effective strategy for mitigating the Rare-Non-rare bias in HOI detection.
Qualitative Results and Analysis
@@
This section presents a qualitative analysis of our method’s predictions, highlighting both its strengths and current limitations. As illustrated in Figures 3(a) and 3(b), our model demonstrates robustness by accurately predicting HOI triplets even when the target object is small or occluded. Conversely, Figure 3(c) presents a representative failure case involving a false positive prediction. Here, the model incorrectly infers that two individuals are driving the boat, a prediction that contradicts common sense in the real world. To diagnose this error, we visualize the corresponding attention maps in Figures 3(d) and 3(e). The visualization reveals that the model correctly identifies promising regions associated with the drive action for both individuals. However, it lacks the commonsense knowledge to understand that only one person can drive a single boat, leading to the erroneous prediction. Furthermore, attention is diffuse and does not pinpoint fine-grained cues, such as hand-on-wheel placement. This limitation highlights a crucial direction for future work: developing methods to inject real-world commonsense reasoning into HOI detectors to resolve such logical ambiguities.
Conclusion
@@
In this paper, we introduce Multimodal Graph Network Modeling (MGNM), a novel and effective two-stage HOI detection framework. The central contribution of MGNM lies in its explicit modeling of the HOI detection task via the construction of a multimodal graph network. Furthermore, by leveraging Vision Language Models (VLMs), we design a multi-level feature interaction mechanism that systematically fuses low-level and high-level vision and language features, thereby enhancing information propagation within the graph structure. Extensive experiments validate that our approach not only achieves state-of-the-art performance but also delivers superior computational efficiency. In addition, our analysis of the pervasive class imbalance issue provides new insights and practical recommendations for mitigating the Rare-Non-rare bias in future research.

### v2

@@ 4.4 Rare-Non-rare Bias
The challenge of few-shot learning is particularly pronounced in HOI detection, where it manifests as a significant performance disparity between rare and non-rare classes. As evidenced in Table 1, this long-tail distribution problem is severe in many methods. For instance, PPDM and FGAHOI exhibit a performance gap of over 10 mAP between their rare and non-rare classes. Similarly, in Figure 4, the GNN-based method SCG also has a significant difference between rare and non-rare classes. However, a key observation from Table 1 is that methods incorporating VLMs substantially mitigate this disparity. This suggests that the rich prior knowledge in VLMs can effectively alleviate the Rare-Non-rare bias. Results illustrated in Figure 4 also support this hypothesis: both MGNM and ViPLO, which leverage CLIP, demonstrate a much narrower performance gap between rare and non-rare classes. Moreover, as shown in Table 1 and Figure 4, an unexpected trend emerges for methods equipped with a more advanced object detector, such as MP-HOI-L, PViC-L, Pose-aware and our $\text{MGNM}_{L}$. These models exhibit a remarkable reversal, where performance on rare classes actually exceeds that on non-rare classes. We posit that this is attributable to the unique architectural design of DDETR [Zhang2022DINODW] or $\mathcal{H}$-DETR with Swin-L backbone, and a deeper investigation into this phenomenon presents a compelling direction for future research. In summary, our analysis indicates that leveraging VLMs and advanced object detectors represents a highly effective strategy for mitigating the Rare-Non-rare bias.
5 Qualitative Results and Limitations
@@
This section presents a qualitative analysis of our method’s predictions, highlighting both its strengths and current limitations. As illustrated in Figures 3(a)-3(e), our model demonstrates robustness by accurately predicting HOI triplets in single-person or multi-person scenarios, even when the target object is small or occluded. Conversely, Figure 3(f) presents a representative failure case involving a false positive prediction. Here, the model incorrectly infers that two individuals are driving the boat, a prediction that contradicts common sense in the real world. To diagnose this error, we visualize the corresponding attention maps in Figures 3(f) and 3(h). The visualization reveals that the model correctly identifies promising regions associated with the drive action for both individuals. However, it lacks common-sense knowledge to understand that only one person can drive a single boat, leading to the erroneous prediction. Furthermore, attention is diffuse and does not pinpoint fine-grained cues, such as hand-on-wheel placement. This limitation highlights a crucial direction for future work: developing methods with real-world common-sense reasoning to resolve such logical ambiguities.
6 Conclusion
@@
In this paper, we introduce multimodal graph network modeling, a novel and effective two-stage HOI detection framework. The central contribution of MGNM lies in its explicit modeling of the HOI detection task via a multimodal graph network. Furthermore, by leveraging vision language models and two designed prompts, we design a multi-level feature interaction mechanism that systematically fuses low-level and high-level visual and language features, thereby enhancing information propagation within the graph structure. Extensive experiments validate that our approach achieves the SOTA performance. In addition, our analysis of the pervasive class imbalance issue provides new insights and practical suggestions for mitigating the Rare-Non-rare bias in future research.